OpenStudy (anonymous):

Given c = 628, b = 480 and C = 55°. The value of a in triangle ABC is 765 and 525 765 525 no solution

OpenStudy (campbell_st):

draw a diagram  this is a 2 part question 1. Find the size of angle B using the sin rule sin(55)/628 = sin(B)/480 sinB = 480 sin(55)/628 \[B = \sin^{-1} (480 \times \sin(55))/628\] use the angle sum of a triangle to find angle A 2. Use the cosine rule to find he length of a \[a^2 = 628^2 + 480^2 - 2 \times 628 \times 480 \times \cos(A)\]